OpenStudy (anonymous):

WILL MEDAL Which statement is true about the equations –3x + 4y = 12 and 1/4x –1/3y = 1? A)The system of the equations has exactly one solution at (–8, 3). B)The system of the equations has exactly one solution at (–4, 3). C)The system of the equations has no solution; the two lines are parallel. D)The system of the equations has an infinite number of solutions represented by either equation.

OpenStudy (anonymous):

PLZ HELP

OpenStudy (anonymous):

do you want answer only or full work?

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

because i dont have time

563blackghost (563blackghost):

You cannot have the answer only....You must understand your question....

563blackghost (563blackghost):

If an openstudier gives away the answer then they would violate the code of conduct...

563blackghost (563blackghost):

To find out if they are parallel you would turn these standard form equations into slope-intercept form...

OpenStudy (anonymous):

either way, here is full text: -3x+4y=12 -3x=12-4y x=-4+4/3y

563blackghost (563blackghost):

For A and B you can input the points to find out if they equal the totals...

OpenStudy (anonymous):

now you change that x in other equation \[1/4*(4/3*y-4) - 1/3y =1\]

OpenStudy (anonymous):

you will see that you will get no answer, which means that answer is C

OpenStudy (anonymous):

yeah thank you so much and I actually understand

OpenStudy (anonymous):

no problem :)

563blackghost (563blackghost):

Nicely Done @stefancipe

OpenStudy (anonymous):

@563blackghost thank you :)